# Policymaker meetings as heteroscedasticity shifters:

### v1

@@ 3.4 Simultaneous confidence bounds
    Use Algoritm 3 to compute $\{\hat{\theta}^{s}\}_{s=1}^{S}$.
@@
    Estimate standard error for each $h=1,\dots,H$
    $\hat{\sigma}_{h}=(q(f_{h}(\hat{\theta}^{s}),\Phi(1))-q(f_{h}(\hat{\theta}^{s}),1-\Phi(1)))/2$ (41)
@@
    $\xi^{s}=\sup_{h=1,\dots,H}\frac{|f_{h}(\hat{\theta}^{s})-f_{h}(\hat{\theta})|}{\hat{\sigma}_{h}}.$ (42)
@@
    Compute
    $[f_{j}(\hat{\theta})-\hat{\sigma}_{h}CV_{\alpha},f_{h}(\hat{\theta})+\hat{\sigma}_{h}CV_{\alpha}].$ (43)
In large samples, by the delta method, the boostrap distribution of $f_{h}(\hat{\theta}^{s})$ is approximately Gaussian. As a result, the difference between the $q(f_{h}(\hat{\theta}^{s}),\Phi(1))$ and $q(f_{h}(\hat{\theta}^{s}),1-\Phi(1))$ quantiles is converging in probablity to 2 standard deviations of $f_{h}(\hat{\theta})$ as $T,S\to\infty$. Step 1 can be replaced with a conventional bootstrap standard error estimator if the bootstrap distribution tails are not fat. Another advantage of using the s.e. estimator based on the quantiles is that the simultaneous bounds coincide with the $68\%$ pointwise bounds if we set the critical value $CV_{\alpha}=1$.
An alternative use of Algorithm 4 is a test of misspecification. Suppose that we have two competing identification assumptions, for example, heteroscedasticity IV and Cholesky. We can construct a joint confidence set for the vector of differences of impacts under these two schemes, $\hat{b}^{Cholesky}_{.1}-\hat{b}^{HetIV}_{.1}$. Alogrithm 4 can then be used to compute $\sup t$ critical values for a test of the hypothesis $b^{Cholesky}_{.1}-b^{HetIV}_{.1}=0$.
@@
3.5 Smoothed local projections using minimum distance approach.
Throughout the paper, we assume that the underlying process that generates the data follows a finite order VAR($p$) model. This model imposes particular relations on the matrices $C_{h}$, $h=1,\dots,H$ ($C_{0}=I$, $C_{h}=\sum_{\ell=1}^{h}A_{\ell}C_{h-\ell}$ with $A_{\ell}=0$ for $\ell>p$) that can be used to improve efficiency. The backward recursion estimator $\hat{A}^{BR}_{i}$, introduced in Algorithm 1, imposes the VAR structure on $C_{h}$ with $h=1,\dots,p$ but does not use the overidentifying information about $C_{h}$ for $h>p$. One can improve efficincy of the LP estimates using the following miminum distance estimator. Take any positive definite matrix $\hat{W}$ that converges to a positive definite matrix $W$ of dimension $n^{2}\times H_{1}$ and minimize
@@
where the vector of local projection is denoted as $\hat{g}^{LP}=vec(\hat{C}_{1},\dots,\hat{C}_{H_{1}})$ and $g(A_{1},\dots,A_{p})=vec(C_{1}(A_{1},\dots,A_{p}),\dots,C_{H_{1}}(A_{1},\dots,A_{p}))$ is the corresponding recursively computed IRF coefficients for any values of matrices $(A_{1},\dots,A_{p})$. From the theory of the minimum distance estimators (see, for example, Hansen (2022)) the optimal choice of $\hat{W}$ correspond to the inverse of the long run covariance matrix, given in Theorem 4 earlier in this paper. In small samples, however, the HAC estimator given in Theorem 5 may result in non-invertible matrix that needs to be regularized. In this paper, we instead use a diagonal matrix $\hat{W}$ with vector of inverse variances of the components of $\hat{g}^{LP}$ on the diagonal. These variances are estimated using dependent wild bootstrap.
Based on the estimates of minimum distance $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$, we can compute smoothed impulse responses for any horizon $H$, $(\hat{C}^{MD}_{1},\dots,\hat{C}^{MD}_{H})\triangleq C_{1}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),\dots,C_{H}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p})$, using the usual recursive formulas.
Under the correct VAR($p$) specification, the limit of (44) coincides with the true parameters $A_{1},\dots,A_{p}$. Moreover, the estimator $((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p})$ is a smooth transformation of $\hat{\theta}$. So, by the delta method, the dependent wild bootstrap from Theorem 6 remains valid for $\hat{C}^{MD}_{i}$. By implication, we can use it both for point-wise inference and simultaneous inference on $C_{h}$.
@@
The idea of smoothing local projection is not new. Barnichon and Brownlees (2019) imposed a particular flexible form on the LP estimates within a ridge-regression framework without providing the corresponding inference procedure. In contrast to the earlier studies, we use the VAR structure of the IRF. This allows us to improve efficiency under correct specification and use the dependent wild bootstrap as a robust inference procedure.
@@ 4 Simulation experiment based on a DSGE model of US economy
In this section, we would like to illustrate how hetersocedasticity IV identification works together with the dependent wild bootstrap using data created by a dynamic general equilibrium model.
@@
For the data generation process, we will use the well-known model Smets and Wouters (2007) (SW later on), which is a medium-scale New Keynesian model with a wide range of nominal and real rigidities. In particular, the model introduces nominal price and wage rigidity, consumption habits, investment adjustment, and capital utilization costs. At the same time, the Taylor rule for the interest rate assumes that the monetary authority reacts to the deviation of output from the level that would prevail in an economy with flexible prices and wages, which further increases the dimension of the problem, since this requires additionally describing the dynamics of a hypothetical economy with flexible prices.
@@
The state vector in the model has a dimension of about 20 variables. As in the original paper, we introduced 7 structural macroeconomic shocks into the model, including the monetary policy shock. Typically, the model is estimated using the Kalman filter with the observation vector consisting of 7 variables (the number of variables cannot be less than the number of shocks): the log difference of real GDP, real consumption, real investment and the real wage, log hours worked, the log difference of the GDP deflator, and the federal funds rate. The filter is used to estimate the linear state space model that corresponds to the solution of the log-linearized DSGE models. The unobserved state vector follows an SVAR model. Since the observation vector has a smaller dimension than the state vector (7 vs 20), the observation vector is not generally a finite-order vector autoregression. However, in a fairly general case, the vector of observed variables can be represented as a VARMA process, and, assuming invertibility of the MA part, a VAR of finite order can act as an approximation for the corresponding VAR($\infty$).  [^8] The quality of approximation by a finite-order VAR depends on the specific structure of the model, the macroeconomics variables used as observables, and on the persistence of shocks (Ravenna, 2007; Fernandez-Villaverde et al., 2007).
@@
In the experiment of this section, we generate samples of various lengths from the log-linearized SW DSGE model (equivalent to a VARMA model) calibrated using the posterior mean estimates obtained in the (Smets and Wouters, 2007). The only exception was the standard deviation parameter of the monetary policy shock, for which we introduced two regimes described by a Markov chain. We consider a symmetric case where the probability of remaining in one of the regimes is 70%, and the switch to another regime is 30%. Then we introduce an auxiliary instrumental variable $Z_{t}$, which takes the value 0 in the first regime and 1 in the second. The value of the standard deviation of the monetary policy shock, $\sigma_{r,t}$, was set according to the following equation:
@@
where $\sigma_{r,0}$ takes a value of posterior mean estimate for the standard deviation of the monetary policy shock in the Smets and Wouters (2007) paper. This specification also assumes that the standard deviation of the monetary shock increases by a factor of 4 in the second regime.
Figure 1: Theoretical versus Empirical Impulse Response Functions to a Monetary Policy Shock identified through the Conditional Heteroscedasticity, sample length is 160 periods.LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line)
Figure 2: Theoretical versus Empirical Impulse Response Functions to a Monetary Policy Shock identified through the Conditional Heteroscedasticity, sample length is 1600 periods. LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line).
Next, using the simulated dataset of seven observable variables and the instrument $Z_{t}$ we apply our new identification method to extract the structural monetary policy shock and compute the corresponding IRFs. We included in the vector of observed variables in local projection model the following variables: log-deviations from the steady state level for real GDP, household consumption, investment, labor, wage, GDP deflator growth rate, and the gross nominal interest rate. To emphasize that the model can handle both stationary and nonstationary variables, we used GDP deflator index rather than growth rates in the econometric exercise. Thus, the model will use 6 stationary variables and 1 non-stationary variable. The original calibration of the model is based on the quarterly frequency of observation. The model includes 4 lags, which corresponds to a lag of 1 year. $Z_{t}$ is used as an instrument for heteroscedasticity of monetary policy shock.
Finally, we compare the IRFs generated from the SW model (VARMA, theoretical) to those obtained from the estimated VAR model to validate our method’s accuracy and effectiveness in capturing the dynamics prescribed by the SW framework. Figure 1 shows the estimation results for 160 periods (40 years), and Figure 2 shows the results for 1600 periods. In Figure 1, it can be seen that the method can reconstruct the IRF for monetary policy shocks within the approximate VAR(4) model using a typical sample size. The true IRF are well within the simultaneous 68% confidence bounds. As we increase the sample size to 1600 periods (Figure 2), the LP estimates are getting even closer to the true values. This shows that LP estimators based on finite VAR(4) can reasonably well approximate the underlying VARMA model. Moreover, the simultaneous confidence bounds successfully covered the true IRF in all the figures, including the non-stationary GDP deflator index.
@@ 5 Empirical Application: identificaiton of monetary policy shock in US
Identification of a monetary policy shock is a classical application of structural VAR models that goes back to Sims (1980). This application has been extensively studied with many identification schemes being proposed, including the classical Cholesky decomposition (Sims, 1980; Christiano et al., 1999), instrumental variables (for example, Romer and Romer, 1989, 2004; Gertler and Karadi, 2015), sign restrictions (for example, Uhlig, 2005; Gafarov et al., 2018; Antolín-Díaz and Rubio-Ramírez, 2018), high-frequency identification (for example, Faust et al., 2004), and regime-switching models (for example, Sims and Zha, 2006; Lütkepohl and Netšunajev, 2017) to name a few. [^9] For foundational context and a detailed examination of prior empirical efforts we refer reader to seminal works by Christiano et al. (1999), Boivin et al. (2010), Ramey (2016), which provide comprehensive reviews on the subject. Such an abundance of benchmarks provides an opportunity to test the novel identification approach and validate or reject selected existing alternative methods.
@@ 5.1 Data and Model
To compare our identification strategy to those commonly used in the literature, we focus on the reduced-form model studied in seminal papers including, but not limited to Christiano et al. (1999), Uhlig (2005), Antolín-Díaz and Rubio-Ramírez (2018). Our specification includes six variables for the U.S.: real GDP, GDP deflator, commodity price index, total reserves, nonborrowed reserves, and federal funds rate, described in Table 1. The sample covers the period from January 1965 to November 2007, with monthly data similar to Antolín-Díaz and Rubio-Ramírez (2018). [^10] The monthly series for real GDP and the GDP deflator are interpolated from quarterly data using Chow and Lin (1971) methodology. We use the industrial production index for real GDP interpolation and the consumer price index for the GDP deflator. Our specification includes a constant and a polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$. [^11] When we compare results to those from the literature, we mimic number of lags and trends. For example, Christiano et al. (1999) uses constant but no time trend, number of lags 6 and 12. Uhlig (2005), Antolín-Díaz and Rubio-Ramírez (2018) use 12 lags and no constant or deterministic trends.
@@
Variable Source Seasonally Mnemonics/Link Adjusted Real GDP FRED Yes https://fred.stlouisfed.org/series/GDPC1 GDP Deflator FRED Yes https://fred.stlouisfed.org/series/GDPDEF) Total Reserves FRED Yes https://fred.stlouisfed.org/series/TRARR Nonborrowed Reserves FRED Yes https://fred.stlouisfed.org/series/BOGNONBR Fed Funds Rate FRED No https://fred.stlouisfed.org/series/FEDFUNDS Commodity Price Index GFD No Global Financial Data Consumer Price Index FRED Yes https://fred.stlouisfed.org/series/CPIAUCSL Industrial Production Index FRED Yes https://fred.stlouisfed.org/series/INDPRO

### v2

@@ 3.4 Simultaneous confidence bounds
    Use Algorithm 3 to compute $\{\hat{\theta}^{s}\}_{s=1}^{S}$.
@@
    Estimate standard error for each $h=1,\dots,H_{2}$
    $\hat{\sigma}_{h}=(q(f_{h}(\hat{\theta}^{s}),\Phi(1))-q(f_{h}(\hat{\theta}^{s}),1-\Phi(1)))/2,$ (41)
@@
    $\xi^{s}=\sup_{h=1,\dots,H_{2}}\frac{|f_{h}(\hat{\theta}^{s})-f_{h}(\hat{\theta})|}{\hat{\sigma}_{h}}.$ (42)
@@
    Compute for all $h$
    $[f_{h}(\hat{\theta})-\hat{\sigma}_{h}CV_{\alpha},f_{h}(\hat{\theta})+\hat{\sigma}_{h}CV_{\alpha}].$ (43)
In large samples, by the delta method, the bootstrap distribution of $f_{h}(\hat{\theta}^{s})$ is approximately Gaussian. As a result, the difference between the $q(f_{h}(\hat{\theta}^{s}),\Phi(1))$ and $q(f_{h}(\hat{\theta}^{s}),1-\Phi(1))$ quantiles is converging in probability to 2 standard deviations of $f_{h}(\hat{\theta})$ as $T,S\to\infty$. Step 1 can be replaced with a conventional bootstrap standard error estimator if the sample size is sufficiently large. [^10] Substitution of interquartile range for standard deviation is frequently used in the bootstrap literature (e.g. Chernozhukov et al., 2020). See, for example, Remark 3.2 of Chernozhukov et al. (2013) which uses the insights from Kato (2011). Another advantage of using the s.e. estimator based on the quantiles is that the simultaneous bounds coincide with the $68\%$ point-wise bounds if we set the critical value $CV_{\alpha}=1$.
An alternative use of Algorithm 4 is a test of misspecification. Suppose that we have two competing identification assumptions, for example, heteroscedasticity IV and Cholesky. We can construct a joint confidence set for the vector of differences of impacts under these two schemes, $\hat{b}^{Cholesky}_{.1}-\hat{b}^{HetIV}_{.1}$. Alogrithm 4 can then be used to compute $\sup t$ critical values for a test of hypothesis $b^{Cholesky}_{.1}-b^{HetIV}_{.1}=0$.
@@
3.5 Smoothed local projections using minimum distance approach
Throughout the paper, we assume that the underlying process that generates the data follows a finite order VAR($p$) model. This model imposes particular relations on the matrices $C_{h}$, $h=1,\dots,H$ that can be used to improve efficiency (i.e., $C_{0}=I$, $C_{h}=\sum_{\ell=1}^{h}A_{\ell}C_{h-\ell}$ with $A_{\ell}=0$ for $\ell>p$). The backward recursion estimator $\hat{A}^{BR}_{i}$, introduced in Algorithm 1, imposes the VAR structure on $C_{h}$ with $h=1,\dots,p$. However, it does not use the overidentifying information about $C_{h}$ for $h>p$. One can improve efficiency of the LP estimates using the following minimum distance estimator. Take any positive definite matrix $\hat{W}$ that converges to a positive definite matrix $W$ of dimension $n^{2}\times H_{1}$ and minimize
@@ 3.5 Smoothed local projections using minimum distance approach
where the vector of the local projection estimators is denoted as $\hat{g}^{LP}=vec(\hat{C}_{1},\dots,\hat{C}_{H_{1}})$ and $g(A_{1},\dots,A_{p})=vec(C_{1}(A_{1},\dots,A_{p}),\dots,C_{H_{1}}(A_{1},\dots,A_{p}))$ is the corresponding recursively computed IRF coefficients for any values of matrices $(A_{1},\dots,A_{p})$. From the theory of the minimum distance estimators (see, for example, Hansen (2022)) the optimal choice of $\hat{W}$ corresponds to the inverse of the long-run covariance matrix, given in Theorem 4 earlier in this paper. In small samples, however, the HAC estimator given in Theorem 5 may result in a non-invertible matrix that needs to be regularized. In this paper, we instead use a diagonal matrix $\hat{W}$ with the vector of inverse variances of the components of $\hat{g}^{LP}$ on the diagonal. These variances are estimated using the dependent wild bootstrap.
Based on the estimates of minimum distance $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$, we can compute smoothed impulse responses for any horizon $H$,
$(\hat{C}^{MD}_{1},\dots,\hat{C}^{MD}_{H})\triangleq C_{1}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),\dots,C_{H}((\hat{A}^{MD}_{1},\dots,(\hat{A}^{MD}_{p}),$
using the usual recursive formulas.
Under the correct VAR($p$) specification, the limit of (44) coincides with the true parameters $A_{1},\dots,A_{p}$. Moreover, the estimator $(\hat{A}^{MD}_{1},\dots,\hat{A}^{MD}_{p})$ is a smooth transformation of $\hat{\theta}$. So, by the delta method, the dependent wild bootstrap from Theorem 6 remains valid for $\hat{C}^{MD}_{i}$. By implication, we can use it both for point-wise inference and simultaneous inference on $C_{h}$.
@@
The idea of smoothing the local projection is not new. Barnichon and Brownlees (2019) imposed a particular flexible form on the LP estimates within a ridge-regression framework without providing the corresponding inference procedure. In contrast to earlier studies, we use the VAR structure of the IRF for smoothing. This allows us to improve efficiency under correct specification and use the dependent wild bootstrap as a robust inference procedure.
@@ 4 Simulation experiment based on a DSGE model of US economy
In this section, we would like to illustrate how heteroscedasticity IV identification works with the dependent wild bootstrap using data created by a dynamic general equilibrium (DSGE) model.
@@
For the data generation process, we will use the well-known model of Smets and Wouters (2007) (SW later on), a medium-scale New Keynesian model with a wide range of nominal and real rigidities. In particular, the model introduces nominal price and wage rigidity, consumption habits, investment adjustment, and capital utilization costs. At the same time, the Taylor rule for the interest rate assumes that the monetary authority reacts to the deviation of output from the level that would prevail in an economy with flexible prices and wages, which further increases the dimension of the problem because this requires additionally describing the dynamics of a hypothetical economy with flexible prices.
@@
The state vector in the model has a dimension of about 20 variables. As in the original paper, we introduce seven structural macroeconomic shocks into the model, including the monetary policy shock. Typically, the model is estimated using the Kalman filter with the observation vector consisting of seven variables (the number of variables cannot be less than the number of shocks): the log difference of real GDP, real consumption, real investment, and the real wage, log hours worked, the log difference of the GDP deflator, and the federal funds rate. The filter is used to estimate the linear state space model that corresponds to the solution of the log-linearized DSGE models. The unobserved state vector follows a SVAR model. Since the observation vector has a smaller dimension than the state vector (7 vs. 20), the observation vector is not generally a finite-order vector autoregression. However, in a fairly general case, the vector of observed variables can be represented as a VARMA process, and, assuming the invertibility of the MA part, a VAR of finite order can act as an approximation for the corresponding VAR($\infty$). [^11] The quality of approximation by a finite-order VAR depends on the specific structure of the model, the macroeconomics variables used as observables, and on the persistence of shocks (Ravenna, 2007; Fernandez-Villaverde et al., 2007).
@@
In the experiment of this section, we generate samples of various lengths from the log-linearized SW DSGE model (equivalent to a VARMA model) calibrated using the posterior mean estimates obtained in (Smets and Wouters, 2007). The only exception was the standard deviation parameter of the monetary policy shock, for which we introduce two regimes described by a Markov chain. We consider a symmetric case where the probability of remaining in one of the regimes is 70%, and the probability of switching to the other regime is 30%. We introduce an auxiliary instrumental variable $Z_{t}$, which takes the value 0 in the first regime and 1 in the second. The value of the standard deviation of the monetary policy shock, $\sigma_{r,t}$, is set according to the following equation:
@@
where $\sigma_{r,0}$ is set to a value of the posterior mean estimate for the standard deviation of the monetary policy shock in the Smets and Wouters (2007) paper. This specification assumes that the standard deviation of the monetary shock increases by a factor of four in the second regime.
Figure 1: Theoretical versus empirical impulse response functions to a monetary policy shock identified through heteroscedasticity-IV, sample length is 160 periods. LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line).
Figure 2: Theoretical versus empirical impulse response functions to a monetary policy shock identified through the heteroscedasticity-IV, sample length is 1600 periods. LP estimator (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line), true IRF from SW model (gold line).
Next, using the simulated dataset of seven observable variables and the instrument $Z_{t}$, we apply our new identification method to extract the structural monetary policy shock and compute the corresponding IRFs. We include in the vector of observed variables in the local projection model the following variables: log deviations from the steady-state level for real GDP, household consumption, investment, labor, wage, GDP deflator index, and the gross nominal interest rate.121212 We use the GDP deflator index rather than growth rates in the econometric exercise in order to emphasize that the method can handle both stationary and nonstationary variables. Thus, the model will use six stationary variables and one nonstationary variable. The original calibration of the model is based on the quarterly frequency of observation. The model includes four lags, which corresponds to a lag of one year. $Z_{t}$ is used as an instrument for the heteroscedasticity of the monetary policy shock.
Finally, we compare the IRFs generated from the theoretical SW model (VARMA) to those obtained from the estimated VAR model to validate our method’s accuracy and effectiveness in capturing the dynamics prescribed by the SW framework. Figure 1 shows the estimation results for 160 periods (40 years), and Figure 2 shows the results for 1600 periods. In Figure 1, it can be seen that the method can reconstruct the IRF for monetary policy shocks within the approximate VAR(4) model using a typical sample size. The true IRFs are well within the simultaneous 68% confidence bounds. As we increase the sample size to 1600 periods (Figure 2), the LP estimates get even closer to the true values. This result shows that LP estimators based on finite VAR(4) can reasonably well approximate the underlying VARMA model. Moreover, the simultaneous confidence bounds successfully covered the true IRFs in all the figures, including the non-stationary GDP deflator index.
@@ 5 Empirical application: identificaiton of monetary policy shock in US
The identification of monetary policy shocks is a classic application of structural VAR models that goes back to Sims (1980). This application has been extensively studied with many identification schemes being proposed, including the classical Cholesky decomposition (Sims, 1980; Christiano et al., 1999), instrumental variables (for example, Romer and Romer, 1989, 2004; Gertler and Karadi, 2015), sign restrictions (for example, Uhlig, 2005; Gafarov et al., 2018; Antolín-Díaz and Rubio-Ramirez, 2018), high-frequency identification (for example, Faust et al., 2004), and regime-switching models (for example, Sims and Zha, 2006; Lütkepohl and Netšunajev, 2017), to name a few. [^13] For foundational context and a detailed examination of prior empirical efforts we refer the reader to seminal works by Christiano et al. (1999), Boivin et al. (2010), Ramey (2016), which provide comprehensive reviews on the subject. Such an abundance of benchmarks provides an opportunity to test the novel identification approach and validate or reject selected existing alternative methods.
@@ 5.1 Data and model
To compare our identification strategy with those commonly used in the literature, we focus on the reduced-form model studied in seminal papers including, but not limited to Christiano et al. (1999), Uhlig (2005), Antolín-Díaz and Rubio-Ramirez (2018). Our specification includes six variables for the U.S.: real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and federal funds rate, as described in Table 1. The sample covers the period from January 1965 to November 2007, with monthly data similar to Antolín-Díaz and Rubio-Ramirez (2018). [^14] The monthly series for real GDP and the GDP deflator are interpolated from quarterly data using the Chow and Lin (1971) methodology. We use the industrial production index for the real GDP interpolation and the consumer price index for the GDP deflator. Our specification includes a constant and a polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$. [^15] When comparing our results to those from the literature, we align the number of lags and the inclusion of trends to match their specifications. For example, Christiano et al. (1999) use constant but without time trend, with lag lengths of 6 and 12, while Uhlig (2005), Antolín-Díaz and Rubio-Ramirez (2018) use 12 lags without a constant or deterministic trend.
@@
Variable Source Seasonally Mnemonics Adjusted Real GDP FRED Yes GDPC1 GDP Deflator FRED Yes GDPDEF) Total Reserves FRED Yes TRARR Nonborrowed Reserves FRED Yes BOGNONBR Federal Funds Rate FRED No FEDFUNDS Commodity Price Index GFD No Global Financial Data Consumer Price Index FRED Yes CPIAUCSL Industrial Production Index FRED Yes INDPRO
Notes: Time series are available on the following website, https://fred.stlouisfed.org/series/.